### v1

@@ Abstract
Transformer-based models have shown strong performance across diverse time-series tasks, but their deployment on resource-constrained devices remains challenging due to high memory and computational demand. While prior work targeting Microcontroller Units (MCUs) has explored hardware-specific optimizations, such approaches are often task-specific and limited to 8-bit fixed-point precision. Field-Programmable Gate Arrays (FPGAs) offer greater flexibility, enabling fine-grained control over data precision and architecture. However, existing FPGA-based deployments of Transformers for time-series analysis typically focus on high-density platforms with manual configuration. This paper presents a unified and fully automated deployment framework for Tiny Transformers on embedded FPGAs. Our framework supports a compact encoder-only Transformer architecture across three representative time-series tasks (forecasting, classification, and anomaly detection). It combines quantization-aware training (down to 4 bits), hardware-aware hyperparameter search using Optuna, and automatic VHDL generation for seamless deployment. We evaluate our framework on six public datasets across two embedded FPGA platforms. Results show that our framework produces integer-only, task-specific Transformer accelerators achieving as low as 0.033 mJ per inference with millisecond latency on AMD Spartan-7, while also providing insights into deployment feasibility on Lattice iCE40. All source code will be released in the GitHub repository [^1] https://github.com/Edwina1030/TinyTransformer4TS.
@@ III-A Overview of Deployment Framework
As shown in Figure 1, our framework begins by designing and configuring a PyTorch-based encoder-only Transformer(❶), referred to as the FloatTransformer, which serves as a unified backbone for three time-series tasks. The model is first trained in Floating-point 32 (FP32) and validated on the target dataset to establish a performance baseline (❷). For quantized deployment, each module in the FloatTransformer is replaced with its quantization-compatible counterpart from an open-source ElasticAI.Creator [^2] https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization library [18], forming the QuantTransformer (❸). QAT is applied to calculate quantization parameters for later enabling integer-only inference. Once training converges, the model is exported to an integer format (❹) suitable for hardware mapping. In the VHDL generation phase (❺), each quantized layer is translated to a RTL module with parameterized bitwidths and tensor dimensions. The design is first functionally verified using GHDL RTL simulation (❻) and then synthesized with Vivado (for AMD Spartan-7) or Radiant (for Lattice iCE40) (❼), generating post-synthesis reports on logic utilization, timing, and estimated power consumption. Inspired by hardware-aware optimization in LOTTA [19], we integrate an Optuna-driven hyperparameter search loop (❽). Each candidate configuration undergoes QAT, simulation, synthesis, and deployability filtering. Valid designs are deployed to real hardware for final validation (❾). We target two embedded FPGA platforms: ElasticNode V5 (AMD Spartan-7 XC7S15 @100 MHz, 8,000 LUTs, 20 DSPs, 10 BRAMs), and ElasticNode V5 SE (Lattice iCE40UP5K @16 MHz, 5,280 LUTs, 8 DSPs, 30 ERBs). Both platforms integrate an RP2040 Cortex-M0+ MCU coordinating data acquisition and managing inference requests to the FPGA [20].
@@
III-B Tiny Transformer Architecture ❶❷
@@
III-C Quantization with ElasticAI.Creator ❸❹
@@
III-D VHDL Generation via Modular Templates ❺
@@
III-E RTL Simulation and FPGA Synthesis ❻❼
@@
III-F Hardware-Aware Optimization ❽
@@ IV-A Experimental Setup
Each trial runs for up to 100 epochs with early stopping (patience = 10). All experiments are conducted in PyTorch 3.11 on an NVIDIA RTX 2080 SUPER GPU (CUDA 11.0). The Adam optimizer is used with default settings. Notably, we fix the operating frequency to 100 MHz for AMD XC7S15 FPGA and 16 MHz for Lattice iCE405UPK FPGA across all experiments for fair comparison.
TABLE I: Energy-efficient Transformer configurations for time-series tasks on AMD XC7S15 FPGA.
@@ IV-B Time-Series Forecasting
Table I summarizes the selected optimal configurations. On the PeMS dataset, the selected model configuration ($d_{\text{model}}\!=\!16$, $bs\!=\!208$, $lr\!=\!6.357\times 10^{-3}$) achieves a test RMSE of 0.194 under 6-bit quantization, representing a 21.25% increase relative to the FP32 baseline (0.160), but a 2.41% improvement over the same-sized 6-bit model reported in [16]. This configuration demonstrates strong deployment efficiency, operating at 1.203 ms latency with only 0.078mJ energy per inference. On the AirU dataset, the selected 8-bit configuration ($d_{\text{model}}\!=\!8$, $bs\!=\!32$) achieves an RMSE of 4.853, 17.53% higher than the FP32 baseline (4.129). Compared to [16], where the same architecture yields an RMSE of 5.377, our model achieves a 9.75% improvement. Furthermore, it surpasses their most energy-efficient variant (5.474 RMSE, 0.084 mJ, 1.33 ms) both in accuracy and efficiency, consuming only 0.036 mJ and 0.570 ms. When benchmarked against the MCU-based deployment in [7] (176 ms latency, 4.048 mJ per inference, 5.44 RMSE), our FPGA-based implementation achieves 308$\times$ lower latency and 112$\times$ lower energy, while also improving accuracy.

### v4

@@ Abstract
Transformer-based models have shown strong performance across diverse time-series tasks, but their deployment on resource-constrained devices remains challenging due to high memory and computational demand. While prior work targeting Microcontroller Units (MCUs) has explored hardware-specific optimizations, such approaches are often task-specific and limited to 8-bit fixed-point precision. Field-Programmable Gate Arrays (FPGAs) offer greater flexibility, enabling fine-grained control over data precision and architecture. However, existing FPGA-based deployments of Transformers for time-series analysis typically focus on high-density platforms with manual configuration. This paper presents a unified and fully automated deployment framework for Tiny Transformers on embedded FPGAs. Our framework supports a compact encoder-only Transformer architecture across three representative time-series tasks (forecasting, classification, and anomaly detection). It combines quantization-aware training (down to 4 bits), hardware-aware hyperparameter search using Optuna, and automatic VHDL generation for seamless deployment. We evaluate our framework on six public datasets across two embedded FPGA platforms. Results show that our framework produces integer-only, task-specific Transformer accelerators achieving as low as 0.033 mJ per inference with millisecond latency on AMD Spartan-7, while also providing insights into deployment feasibility on Lattice iCE40. All source code will be released in the GitHub repository [^1] https://github.com/tianheng-ling/TinyTransformer4TS.
@@ III-A Overview of Deployment Framework
As shown in Figure 1, our framework begins by designing and configuring a QuantTransformer model (❷) using quantization-compatible layer or module from an open-source ElasticAI.Creator [^2] https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization library [18]. QAT is applied to calculate quantization parameters for later enabling integer-only inference. Once training converges, the model is exported to an integer format (❸) suitable for hardware mapping. In the VHDL generation phase (❹), each quantized layer is translated to a RTL module with parameterized bitwidths and tensor dimensions. The design is first functionally verified using GHDL RTL simulation (❺) and then synthesized with Vivado (for AMD Spartan-7) or Radiant (for Lattice iCE40) (❻), generating post-synthesis reports on logic utilization, timing, and estimated power consumption. Inspired by hardware-aware optimization in LOTTA [19], we integrate an Optuna-driven hyperparameter search loop (❼). Each candidate configuration undergoes QAT, simulation, synthesis, and deployability filtering. Only configurations that pass hardware constraints proceed to final validation. To establish a performance baseline, each selected configuration is also evaluated using a PyTorch-based FloatTransformer model (❽) trained and validated on the same dataset in full-precision (FP32). The final quantized models are then deployed on real hardware for end-to-end testing and performance validation (❾). We target two embedded FPGA platforms: ElasticNode V5 (AMD Spartan-7 XC7S15 @100 MHz, 8,000 LUTs, 20 DSPs, 10 BRAMs), and ElasticNode V5 SE (Lattice iCE40UP5K @16 MHz, 5,280 LUTs, 8 DSPs, 30 ERBs). Both platforms integrate an RP2040 Cortex-M0+ MCU coordinating data acquisition and managing inference requests to the FPGA [20].
@@
III-B Tiny Transformer Architecture ❶❽
@@
III-C Quantization with ElasticAI.Creator ❷❸
@@
III-D VHDL Generation via Modular Templates ❹
@@
III-E RTL Simulation and FPGA Synthesis ❺❻
@@
III-F Hardware-Aware Optimization ❼
@@ IV-A Experimental Setup
Each trial runs for up to 100 epochs with early stopping (patience = 10). All experiments are conducted in Python 3.11 on an NVIDIA RTX 2080 SUPER GPU (CUDA 11.0). The Adam optimizer is used with default settings. Notably, we fix the operating frequency to 100 MHz for AMD XC7S15 FPGA and 16 MHz for Lattice iCE405UPK FPGA across all experiments for fair comparison.
TABLE I: Selected Transformer configurations for time-series tasks on AMD XC7S15 FPGA.
@@ IV-B Time-Series Forecasting
Table I lists the Pareto-optimal configurations with the highest performance under integer-only quantization. On the PeMS dataset, the selected model configuration ($d_{\text{model}}\!=\!16$, $bs\!=\!208$, $lr\!=\!6.357\times 10^{-3}$) achieves a test RMSE of 0.194 under 6-bit quantization, representing a 21.25% increase relative to the FP32 baseline (0.160), but a 2.41% improvement over the same-sized 6-bit model reported in [16]. This configuration demonstrates strong deployment efficiency, operating at 1.203 ms latency with only 0.078mJ energy per inference. On the AirU dataset, the selected 8-bit configuration ($d_{\text{model}}\!=\!8$, $bs\!=\!32$) achieves an RMSE of 4.853, 17.53% higher than the FP32 baseline (4.129). Compared to [16], where the same architecture yields an RMSE of 5.377, our model achieves a 9.75% improvement. Furthermore, it surpasses their most energy-efficient variant (5.474 RMSE, 0.084 mJ, 1.33 ms) both in accuracy and efficiency, consuming only 0.036 mJ and 0.570 ms. When benchmarked against the MCU-based deployment in [7] (176 ms latency, 4.048 mJ per inference, 5.44 RMSE), our FPGA-based implementation achieves 308$\times$ lower latency and 112$\times$ lower energy, while also improving accuracy.